# Distributed Solving of Linear Quadratic Optimal Controller with Terminal State Constraint

### v3

@@ 4 Main results
    Let $u_{i}^{*}(t)=-R^{-1/2}W_{i,\varpi}^{\prime}\big{[}P_{i,k}^{n}(t)x_{i,w}^{*}(t)+\beta_{i}(t)\big{]}$ and output $u_{i}^{*}(t)$ and $x_{i,w}^{*}(t)$.
Based on Algorithm 2, the distributed design of the optimal controller (58) is shown in Algorithm 3, the convergence of which is given below.
Algorithm 3 Distributed Algorithm without Terminal State Constraint
Input: Let the initial matrices be $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{\bar{n}}(t)=0$, $P_{i,\bar{n}}^{0}(t)=0$, $W_{i,0}=0$, and $x_{i,0}^{\diamond}(t)=0$ and select a sufficient small constant $\bar{\varsigma}>0$.
Output: The distributed controller $u_{i}^{\diamond}(t)$, $i=1,2,\ldots,N$.
  1. 1.
    For $\bar{n}=1,2,\cdots$,
    • For $\bar{k}=1,2,3,\cdots$, run the iteration (31)-(32) and obtain $Z_{i,\bar{k}}^{\bar{n}}(t)$, $V_{i,\bar{k}}^{\bar{n}}(t)$. If $\|Z_{i,\bar{k}}^{n}(t)-Z_{i,\bar{k}-1}^{\bar{n}}(t)\|<\bar{\varsigma}$ and $\|V_{i,\bar{k}}^{n}(t)-V_{i,\bar{k}-1}^{\bar{n}}(t)\|<\bar{\varsigma}$, end the iteration, and derive $Z_{i,\bar{k}}^{\bar{n}}(t)$ and $V_{i,\bar{k}}^{\bar{n}}(t)$.
    • Solve the equation (30) and obtain $P_{i,\bar{k}}^{\bar{n}}(t)$. If $\|P_{i,\bar{k}}^{\bar{n}}(t)-P_{i,\bar{k}}^{\bar{n}-1}(t)\|<\bar{\varsigma}$, end the iteration, and derive $P^{\bar{n}}_{i,\bar{k}}(t)$, $\Phi^{\bar{n}}_{i,\bar{k}}(t)$.
  2. 2.
    For $\bar{\varpi}=1,2,\cdots$, run the iteration
    $W_{i,\bar{\varpi}}$ $=$ $W_{i,\bar{\varpi}-1}+\alpha_{\bar{\varpi}}[R^{-\frac{1}{2}}B_{i}^{\prime}-W_{i,\bar{\varpi}-1}]$
    $+\frac{1}{\gamma}\sum_{j\in N_{i}}[W_{j,\bar{\varpi}-1}-W_{i,\bar{\varpi}-1}].$
    • If $\|W_{i,\bar{\varpi}}-W_{i,\bar{\varpi}-1}\|<\bar{\varsigma}$, end the iteration, and derive $W_{i,\bar{\varpi}}$.
  3. 3.
    For $\bar{w}=1,2,\cdots$, run the iteration
    $x_{i,\bar{w}}^{\diamond}(t)$ $=$ $x_{i,\bar{w}-1}^{\diamond}(t)+\alpha_{\bar{w}}[\Phi_{i,\bar{k}}^{n}(t,0)x_{i}(0)-x_{i,\bar{w}-1}^{\diamond}(t)]$
    $+\frac{1}{\gamma}\sum_{j\in N_{i}}[x_{j,\bar{w}-1}^{\diamond}(t)-x_{i,\bar{w}-1}^{\diamond}(t)].$
    • If $\|x_{i,\bar{w}}^{\diamond}(t)-x_{i,\bar{w}-1}^{\diamond}(t)\|<\bar{\varsigma}$, end the iteration, and derive $x_{i,\bar{w}}^{\diamond}(t)$.
  4. 4.
    Let $u_{i}^{\diamond}(t)=-R^{-1/2}W_{i,\bar{\varpi}}^{\prime}P_{i,\bar{k}}^{\bar{n}}(t)x_{i,\bar{w}}^{\diamond}(t)$ and output $u_{i}^{\diamond}(t)$ and $x_{i,\bar{w}}^{\diamond}(t)$.
Corollary 4.6.
There exist integers $\bar{n},\bar{k},\bar{\varpi},\bar{w}$ and sufficiently small constants $\bar{\psi}_{1}>0$, $\bar{\psi}_{2}>0$ such that output $x_{i,\bar{w}}^{\diamond}(t)$ and $u_{i}^{\diamond}(t)$ derived by Algorithm 3 satisfies
$\|x_{i,\bar{w}}^{\diamond}(t)-x^{\diamond}(t)\|\leq\bar{\psi}_{1},$ (60)
$\|u_{i}^{\diamond}(t)-u^{\diamond}(t)\|\leq\bar{\psi}_{2},$ (61)
where $x^{\diamond}(t)$ and $u^{\diamond}(t)$ are the optimal solution of (59) and (58) to the standard LQ problem without terminal state constraint by assuming all the information of system parameters are available to make decision.
Proof. The proof is similar to that of Theorem 3 and is thus omitted here.
Remark 4.7.
The distributed solving of the finite-horizon solution (58) has been given in Algorithms 3. When the horizon $T$ is sufficiently large, the output of Algorithm 3
$u_{i}^{\diamond}(t)=-R^{-1/2}W_{i,\bar{\varpi}}^{\prime}P_{i,\bar{k}}^{\bar{n}}(T)x_{i,\bar{w}}^{\diamond}(t).$ (62)
is an approximation solution to the standard infinite-horizon LQ problem under the assumptions of the stabilization of $(A,B)$ and observability of $(A,Q)$.
5 Simulation Example
@@
Example 1. Consider the distributed optimal control of a heterogeneous UGVs system, where the dynamics of each UGV is given by [35]
@@
while $q_{i}(t)\in R^{2}$, $i=1,\cdots,5$, denote the position and velocity vectors of the $i$th UGV, $F_{i}(t)\in R^{2}$ denotes the force applied to the UGV, $D_{i}$ denotes the mechanical mass and $C_{i}$ denotes the translational friction coefficient. The communication topology among the UGVs is illustrated in Figure 2.
Figure 2. Communication topology among heterogeneous UGVs.
@@
By defining the state variables as $x(t)=[x_{1}(t)^{\prime}~{}\cdots~{}x_{5}(t)^{\prime}]^{\prime}$ and the input variable as $u(t)=[u_{1}(t)^{\prime}~{}\cdots~{}u_{5}(t)^{\prime}]^{\prime}$, where ${x}_{i}(t)=[q_{i}(t)^{\prime}~{}~{}v_{i}(t)^{\prime}]^{\prime}$, ${u}_{i}(t)=F_{i}(t)$, then the UGV system (63) can be formulated as (3) with
$A_{i}=\left[\begin{array}[]{cccc}0{}&{}\hfil&{}\hfil&\\ {}\hfil&5\tilde{A}_{i}{}&{}\hfil&\\ {}\hfil&{}\hfil&\ddots{}&\\ {}\hfil&{}\hfil&{}\hfil&0\\ \end{array}\right],B_{i}=\left[\begin{array}[]{cccc}0{}&{}\hfil&{}\hfil&\\ {}\hfil&5\tilde{B}_{i}{}&{}\hfil&\\ {}\hfil&{}\hfil&\ddots{}&\\ {}\hfil&{}\hfil&{}\hfil&0\\ \end{array}\right],$
$\tilde{A}_{i}=\left[\begin{array}[]{cc}0&I\\ 0&-\frac{C_{i}}{D_{i}}I\end{array}\right],~{}~{}\tilde{B}_{i}=\left[\begin{array}[]{c}0\\ \frac{1}{D_{i}}I\end{array}\right].$
The control objective is to drive the UGVs from initial positions to predefined target formation while minimizing the cost functions (2) with $Q=16I$, $R=10I$, the initial states of the UGVs being
$x_{1}(0)=\left[\begin{array}[]{cccc}1.5\\ 2.5\\ 0\\ 0\end{array}\right],x_{2}(0)=\left[\begin{array}[]{cccc}2.5\\ 2.6\\ 0\\ 0\end{array}\right],x_{3}(0)=\left[\begin{array}[]{cccc}2\\ 1\\ 0\\ 0\end{array}\right],$
$x_{4}(0)=\left[\begin{array}[]{cccc}0\\ 1\\ 0\\ 0\end{array}\right],x_{5}(0)=\left[\begin{array}[]{cccc}1\\ 2\\ 0\\ 0\end{array}\right].$
and the terminal states of the UGVs being
$x_{1}(T)=\left[\begin{array}[]{cccc}5.5\\ 8\\ 3\\ 4\end{array}\right],x_{2}(T)=\left[\begin{array}[]{cc}7\\ 7\\ 2\\ 4\end{array}\right],x_{3}(T)=\left[\begin{array}[]{cc}8\\ 6\\ 3\\ 6\end{array}\right],$
$x_{4}(T)=\left[\begin{array}[]{cc}3\\ 6\\ 1\\ 3\end{array}\right],x_{5}(T)=\left[\begin{array}[]{cc}4\\ 7\\ 2\\ 1\end{array}\right].$
Next, by running Algorithm 2 with step size $\alpha_{k}=\frac{1}{k}$ and $\gamma=0.4$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figure 3-Figure 8. It is obvious that all the agents achieve consensus and converge to the centralized solutions from Figure 7-Figure 8. In particular, the trajectories of the agent achieve the predefined target formation at the given time $T=2$s as shown in Figure 9.
Figure 3. The trajectories of $\|P_{i,k}^{n}(t)\|_{2}$.
Figure 4. The trajectories of $\|W_{i,\varpi}\|_{2}$.
Figure 5. The trajectories of $\|\lambda_{i,q}^{*}\|_{2}$.
Figure 6. The trajectories of $\|\beta_{i}(t)\|_{2}$.
Figure 7. The trajectories of $\|x_{i,w}^{*}(t)\|_{2}$ with $w=200$.
Figure 8. The trajectories of $u_{i}^{*}(t)$.
Figure 9. The predefined target formation of heterogeneous UGVs.
$A=\left[\begin{array}[]{cc}1{}&2\\ 0{}&1\end{array}\right],B=\left[\begin{array}[]{cc}1\\ 1\end{array}\right],Q=I,R=1,$ (74)
Moreover, 4 agents are involved in the sensor network and the topology graph is shown in Figure 10.
Figure 10. Communication topology among agents.
$A_{1}=\left[\begin{array}[]{cc}1.5{}&2\\ 0{}&0.5\end{array}\right],A_{2}=\left[\begin{array}[]{cc}0.5{}&2\\ 0{}&1.2\end{array}\right],A_{3}=\left[\begin{array}[]{cc}1.3{}&2\\ 0{}&1.8\end{array}\right],$
$A_{4}=\left[\begin{array}[]{cc}0.7{}&2\\ 0{}&0.5\end{array}\right],B_{1}=\left[\begin{array}[]{cc}1.5\\ 0.5\end{array}\right],B_{2}=\left[\begin{array}[]{cc}0.5\\ 1.2\end{array}\right],$
$B_{3}=\left[\begin{array}[]{cc}0.3\\ 1.8\end{array}\right],B_{4}=\left[\begin{array}[]{cc}1.7\\ 0.5\end{array}\right],Q_{1}=\left[\begin{array}[]{cc}1.2{}&0\\ 0{}&1\end{array}\right],$
$Q_{2}=\left[\begin{array}[]{cc}1{}&0\\ 0{}&0.8\end{array}\right],Q_{3}=\left[\begin{array}[]{cc}1{}&0\\ 0{}&1.2\end{array}\right],Q_{4}=\left[\begin{array}[]{cc}0.8{}&0\\ 0{}&1\end{array}\right],$
$R=1,x_{1,0}=\left[\begin{array}[]{cc}3.5\\ 4\end{array}\right],x_{2,0}=\left[\begin{array}[]{cc}4.5\\ 3.5\end{array}\right],x_{3,0}=\left[\begin{array}[]{cc}2.5\\ 4.5\end{array}\right],$
$x_{4,0}=\left[\begin{array}[]{cc}5.5\\ 4\end{array}\right],x_{i,T}=\left[\begin{array}[]{cc}0\\ 0\end{array}\right],i=1,\cdots,4.$
By running Algorithm 2 with $\alpha_{k}=\frac{1}{k}$, $\gamma=0.4$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figure 11-Figure 16. It is obvious that the distributed solutions converge to the centralized solutions Figure 15-Figure 16.
Figure 11. The trajectories of $\|P_{i,k}^{n}(t)\|_{2}$.
Figure 12. The trajectories of $\|W_{i,\varpi}\|_{2}$.
Figure 13. The trajectories of $\|\lambda_{i,q}^{*}\|_{2}$.
Figure 14. The trajectories of $\|\beta_{i}(t)\|_{2}$.
Figure 15. The trajectories of $\|x_{i,w}^{*}(t)\|_{2}$.
Figure 16. The trajectories of $u_{i}^{*}(t)$.
6 Conclusions

### v4

@@ IV Main results
    Let $u_{i}^{*}(t)=-R^{-1/2}W_{i,\varpi}\big[P_{i,k}^{n}(t)x_{i,w}^{*}(t)+\beta_{i}(t)\big]$ and output $u_{i}^{*}(t)$ and $x_{i,w}^{*}(t)$.
Theorem 3
$\|x_{i,w}^{*}(t)-x^{*}(t)\|\leq\psi_{1},$ (46)
$\|u_{i}^{*}(t)-u^{*}(t)\|\leq\psi_{2},$ (47)
Proof. From Lemmas 1-2 and Corollary 2, there always exists integer $n$, $k$ and a sufficient small constant $\sigma_{1}>0$ such that $\|P_{i,k}^{n}(t)-P(t)\|\leq\sigma_{1},t\in[0,T]$. Together with Lemmas 3-4, this further implies that there exist iteration numbers $\varpi,q$ and sufficient small constants $\sigma_{i}>0,i=2,\cdots,4$ such that
$\|W_{i,\varpi}-W\|\leq\sigma_{2},$
$\|\lambda_{i,q}^{*}-\lambda^{*}\|\leq\sigma_{3},$
$\|\beta_{i}(t)-\beta(t)\|\leq\sigma_{4}.$
Accordingly, by using Lemma 5, there exist iteration numbers $w$ and sufficient small constants $\psi_{1}>0$, $\psi_{2}>0$ such that the outputs $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$ obtained by running Algorithm 2 satisfies (49) and (47). This completes the proof. $\blacksquare$
V Application to the Optimal Consensus Control Problem
In this section, we will utilize Algorithm 2 to solve the consensus problem in multi-agent systems, which has broad applications in fields such as unmanned aerial vehicle formation, sensor networks, smart grid dispatch, and distributed robotic collaboration. Specifically, consider the following heterogeneous multi-agent system consisting of $N$ agents with dynamics described by:
$\dot{x}_{i}(t)=A_{i}x_{i}(t)+B_{i}u_{i}(t),$ (48)
where $x_{i}(t)\in R^{n}$, $u_{i}(t)\in R^{m}$, $i=1,..,N$. The initial state is given by $x_{i}(0)=x_{i,0}$.
To apply Algorithm 2 and achieve consensus, we define the following performance index:
$J$ $=\int_{0}^{\infty}\Big[\sum_{i=1}^{N}\sum_{j\in N_{i}}(x_{i}(t)-x_{j}(t))^{\prime}Q_{ij}(x_{i}(t)-x_{j}(t))$
$\quad\quad\quad~~+u^{\prime}(t)Ru(t)\Big]dt$
$=\int_{0}^{\infty}[x^{\prime}(t)\tilde{Q}x(t)+u^{\prime}(t)Ru(t)]dt,$ (49)
where $x(t)=[x_{1}^{\prime}(t)~\cdots~x_{N}^{\prime}(t)]^{\prime}$, $u(t)=[u_{1}^{\prime}(t)~\cdots~u_{N}^{\prime}(t)]^{\prime}$, $R=diag\{R_{1},R_{2},\ldots,R_{N}\}$ is a positive-definite matrix with compatible dimension, $Q_{ij}$, $j\in N_{i}$ is a positive definite matrix with compatible dimension, and
$\tilde{Q}$ $=\begin{bmatrix}\tilde{Q}_{11}&\tilde{Q}_{12}&\cdots&\tilde{Q}_{1N}\\ \vdots&\vdots&\cdots&\vdots\\ \tilde{Q}_{N1}&\tilde{Q}_{N2}&\cdots&\tilde{Q}_{NN}\end{bmatrix}$,
with
$\tilde{Q}_{ii}$ $=\sum_{j=1}^{N}c_{ij}(Q_{ij}+Q_{ji}),$
$\tilde{Q}_{ij}$ $=-c_{ij}(Q_{ij}+Q_{ji}),$
while $c_{ij}=1$ if $j\in N_{i}$, otherwise $c_{ij}=0$. In particular, each agent $i$ has access to the $i$-th block row of $\tilde{Q}$, i.e., $\tilde{Q}_{i}=L_{i}\tilde{Q}$ with $L_{i}=diag\{0,\cdots,0,NI,0,\cdots,0\}$.
The objective of this section is to design a distributed controller by using its own and neighboring information of the system (48) to minimize the cost function (49).
Remark 3
In the literature, a commonly used protocol for multi-agent systems [37, 38, 39] is:
$u_{i}(t)=K\sum_{j\in N_{i}}(x_{j}(t)-x_{i}(t)),$ (50)
which ensures that the system (48) achieves state consensus, that is, for any initial state $x_{i}(0)$,
$\lim_{t\rightarrow\infty}\|x_{i}(t)-x_{j}(t)\|=0,~~i,j=1,2,\cdots,N.$ (51)
However, such protocols do not consider the optimality, and the optimality of some specific performance index, which is essential in practical scenarios such as smart grids [16] and unmanned ground vehicle (UGV) swarm coordination [32]. In this section, we present an optimal consensus protocol using Algorithm 2.
To address this, we first present the centralized optimal controller, which requires information from all agents.
Lemma 6
Assume that $(A_{i},B_{i})$ is stabilizable and $(A,\tilde{Q}^{1/2})$ is observable, where $A=diag\{A_{1},\cdots,A_{N}\}$. Then, the centralized optimal controller is given by
$u^{\diamond}(t)$ $=-R^{-1}B^{\prime}Px^{\diamond}(t),$ (52)
where $B=diag\{B_{1},\cdots,B_{N}\}$, $P$ is the unique positive-definite solution to the algebraic Riccati equation:
$A^{\prime}P+PA+\tilde{Q}-PBR^{-1}B^{\prime}P$ $=0,$ (53)
and $x^{\diamond}(t)$ satisfies:
$\dot{x}^{\diamond}(t)$ $=[A-BR^{-1}B^{\prime}P]x^{\diamond}(t),\quad x^{\diamond}(0)=x(0).$ (54)
Proof. Since $(A_{i},B_{i})$ is stabilizable, the block diagonal $(A,B)$ is stabilizable. Combined with the observability of $(A,\tilde{Q}^{1/2})$, the algebraic Riccati equation (53) admits a unique positive-definite solution $P$. The optimal controller (52) and the corresponding closed-loop system (54) then follow directly from standard LQR theory [34]. This completes the proof. $\blacksquare$
Next, we derive the distributed solution by using Algorithm 2. To this end, we define
$\tilde{A}_{i}=diag\{0,\cdots,0,NA_{i},0,\cdots,0\},$
$\tilde{B}_{i}=diag\{0,\cdots,0,NB_{i},0,\cdots,0\},$
$\tilde{R}_{i}=diag\{0,\cdots,(1/N)R_{i}^{-1},\cdots,0\},$
and introduce the distributed iteration equation for $P$:
$0$ $=(\bar{Z}_{i,k}^{n})^{\prime}P_{i,k}^{n}+P_{i,k}^{n}\bar{Z}_{i,k}^{n}+\bar{V}_{i,k}^{n},$ (55)
$\bar{Z}_{i,k}^{n}$ $=\bar{Z}_{i,k-1}^{n}+\alpha_{k}[\tilde{A}_{i}-\tilde{B}_{i}\tilde{R}_{i}\tilde{B}_{i}^{\prime}P_{i,k-1}^{n-1}-\bar{Z}_{i,k-1}^{n}]$
$\quad+\frac{1}{\gamma}\sum_{j\in N_{i}}[\bar{Z}_{j,k-1}^{n}-\bar{Z}_{i,k-1}^{n}],$ (56)
$\bar{V}_{i,k}^{n}$ $=\bar{V}_{i,k-1}^{n}+\alpha_{k}[\tilde{Q}_{i}+P_{i,k-1}^{n-1}\tilde{B}_{i}\tilde{R}_{i}\tilde{B}_{i}^{\prime}P_{i,k-1}^{n-1}$
$\quad-\bar{V}_{i,k-1}^{n}]+\frac{1}{\gamma}\sum_{j\in N_{i}}[\bar{V}_{j,k-1}^{n}-\bar{V}_{i,k-1}^{n}],$ (57)
with $\bar{Z}_{i,0}^{n}=0$, $P_{i,k}^{0}=0$, $P_{i,0}^{n}=0$, $\bar{V}_{i,0}^{n}=0$, and the distributed iteration equation for $x^{\diamond}(t)$:
$x_{i,k}^{\diamond}(t)$ $=x_{i,k-1}^{\diamond}(t)+\alpha_{k}\big[e^{\bar{Z}_{i,\infty}t}\bar{x}_{i}(0)-x_{i,k-1}^{\diamond}(t)\big]$
$\quad+\frac{1}{\gamma}\sum_{j\in\mathcal{N}_{i}}[x_{j,k-1}^{\diamond}(t)-x_{i,k-1}^{\diamond}(t)],$ (58)
with $\bar{x}_{i}(0)=[0~\cdots~0~Nx_{i,0}^{\prime}~0~\cdots~0]^{\prime}$, $x_{i,0}^{\diamond}(t)=0$ and $\bar{Z}_{i,\infty}$ is the convergent value of (56).
Then, we present the distributed design of the controller (52) as follows.
Theorem 4
Under the stabilizability of $(A_{i},B_{i})$ and observability of $(A,\tilde{Q}^{1/2})$, let $\alpha_{k}$ satisfy $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, $\sum_{k=0}^{\infty}\alpha_{k}^{2}<\infty$, and choose $\gamma>0$ such that $I_{N}-\frac{1}{\gamma}\mathcal{L}-\frac{1}{N}\mathbf{1}_{N}\mathbf{1}_{N}^{\prime}$ is Hurwitz, then it holds that:
$P_{i,\infty}$ $\triangleq\lim_{n\to\infty}\lim_{k\to\infty}P_{i,k}^{n}=P,$~~
$\bar{Z}_{i,\infty}$ $\triangleq\lim_{n\to\infty}\lim_{k\to\infty}\bar{Z}_{i,k}^{n}=A-BR^{-1}B^{\prime}P,$
$x_{i,\infty}^{\diamond}(t)$ $\triangleq\lim_{k\to\infty}x_{i,k}^{\diamond}(t)=x^{\diamond}(t).$
In this case, the distributed optimal controller for each agent $i$ is given by
$u_{i}(t)=-[0~\ldots~0~R_{i}^{-1}B_{i}^{\prime}~0~\ldots~0]P_{i,\infty}x_{i,\infty}^{\diamond}(t),$ (59)
which satisfies
$\|u(t)-u^{\diamond}(t)\|\leq\varepsilon,$ (60)
and ensures that the closed-loop system (48) achieves consensus.
Proof. The proof is divided into two parts. First, we prove that (60) holds. Similar to the proof of Lemma 2, there exist integers $\bar{n}\geq n,\bar{k}\geq k$ and sufficiently small constants $\varepsilon_{1}>0$, $\varepsilon_{2}>0$ such that
$\|P_{i,\bar{k}}^{\bar{n}}-P\|$ $\leq\varepsilon_{1},$ (61)
$\|x_{i,\bar{k}}^{\diamond}(t)-x^{\diamond}(t)\|$ $\leq\varepsilon_{2}.$ (62)
This implies that the distributed solutions $P_{i,k}^{n}$ and $x_{i,k}^{\diamond}(t)$ can approximate their global solution $P$ and $x^{\diamond}(t)$. Then, combining the fact that $u(t)=[u_{1}^{\prime}(t)~\cdots~u_{N}^{\prime}(t)]^{\prime}$ and
$\begin{bmatrix}R_{1}^{-1}B_{1}^{\prime}&0&\cdots&0\\ 0&R_{2}^{-1}B_{2}^{\prime}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&R_{N}^{-1}B_{N}^{\prime}\end{bmatrix}=R^{-1}B^{\prime},$
we can design the distributed controller (59) from (52) and obtain (60) holds.
Second, we show that the system achieves consensus. Since the problem has a solution and $J<\infty$, then it follows from the properties of infinite integrals that
$\lim_{t\rightarrow\infty}(x_{i}(t)-x_{j}(t))^{\prime}Q_{ij}(x_{i}(t)-x_{j}(t))=0.$ (63)
Combined with $Q_{ij}>0$, the necessary condition for the term (63) to converge to zero over the infinite horizon is
$\lim_{t\rightarrow\infty}\|x_{i}(t)-x_{j}(t)\|=0,~~i,j=1,2,\cdots,N.$
This completes the proof. $\blacksquare$
VI Simulation Example
@@
$A=\left[\begin{array}[]{cc}1&0\\ 0&1\end{array}\right],B=\left[\begin{array}[]{cc}1\\ 1\end{array}\right],Q=I,R=1,$ (68)
Moreover, 4 agents are involved in the sensor network and the topology graph is shown in Fig. 1.
Fig. 1. Communication topology among agents.
$A_{1}=\left[\begin{array}[]{cc}1.5&0\\ 0&0.5\end{array}\right],A_{2}=\left[\begin{array}[]{cc}0.5&0\\ 0&1.2\end{array}\right],$
$A_{3}=\left[\begin{array}[]{cc}1.3&0\\ 0&1.8\end{array}\right],A_{4}=\left[\begin{array}[]{cc}0.7&0\\ 0&0.5\end{array}\right],$
$B_{1}=\left[\begin{array}[]{cc}1.5\\ 0.5\end{array}\right],B_{2}=\left[\begin{array}[]{cc}0.5\\ 1.2\end{array}\right],B_{3}=\left[\begin{array}[]{cc}0.3\\ 1.8\end{array}\right],$
$B_{4}=\left[\begin{array}[]{cc}1.7\\ 0.5\end{array}\right],Q_{1}=\left[\begin{array}[]{cc}1.2&0\\ 0&1\end{array}\right],Q_{2}=\left[\begin{array}[]{cc}1&0\\ 0&0.8\end{array}\right],$
$Q_{3}=\left[\begin{array}[]{cc}1&0\\ 0&1.2\end{array}\right],Q_{4}=\left[\begin{array}[]{cc}0.8&0\\ 0&1\end{array}\right],R=1,$
$x_{1,0}=\left[\begin{array}[]{cc}3.5\\ 4\end{array}\right],x_{2,0}=\left[\begin{array}[]{cc}4.5\\ 3.5\end{array}\right],x_{3,0}=\left[\begin{array}[]{cc}2.5\\ 4.5\end{array}\right],$
$x_{4,0}=\left[\begin{array}[]{cc}5.5\\ 4\end{array}\right],x_{i,T}=\left[\begin{array}[]{cc}0\\ 0\end{array}\right],M_{i}=\left[\begin{array}[]{cc}1&0\\ 0&1\end{array}\right],$
$i=1,\cdots,4.$
By running Algorithm 2 with $\alpha_{k}=\frac{1}{k}$, $\gamma=2.5$, $\varsigma=10^{-3}$, $n=20$, $k=\varpi=q=w=200$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figs. 2-7. It is obvious that the distributed solutions converge to the centralized solutions from the Figs. 2-7.
Fig. 2. The trajectories of $\|P_{i,k}^{n}(t)\|_{2}$.
Fig. 3. The trajectories of $\|W_{i,\varpi}\|_{2}$.
Fig. 4. The trajectories of $\|\lambda_{i,q}^{*}\|_{2}$.
Fig. 5. The trajectories of $\|\beta_{i}(t)\|_{2}$.
Fig. 6. The trajectories of $\|x_{i,w}^{*}(t)\|_{2}$.
Fig. 7. The trajectories of $u_{i}^{*}(t)$.
Example 2. In the example, we will apply Theorem 4 to solve the consensus problem of a heterogeneous UGV system. Specifically, the objective is to enable five UGVs to achieve position consensus, where the dynamics of each UGV are given by [35]:
@@
while $q_{i}(t)\in R^{2}$, $v_{i}(t)\in R^{2}$, $i=1,\cdots,5$, denote the position and velocity vectors of the $i$th UGV, $F_{i}(t)\in R^{2}$ denotes the force applied to the UGV, $D_{i}$ denotes the mechanical mass and $C_{i}$ denotes the translational friction coefficient. The system parameters are listed in Table I, and the communication topology among UGVs is illustrated in Fig. 8.
Fig. 8. Communication topology among heterogeneous UGVs.
TABLE I: Parameters of the UGV system.
Parameters UGV1 UGV2 UGV3 UGV4 UGV5
$C_{i}$ $0.6$ $0.8$ $1.2$ 11 $0.4$
$D_{i}$ 55 44 66 44 33
@@
By defining the state variables as $x_{i}(t)=[q_{i}(t)^{\prime}~~v_{i}(t)^{\prime}]^{\prime}$ and the control input as $u_{i}(t)=F_{i}(t)$, then the UGV system (81) can be formulated as (48) with
$A_{i}=\left[\begin{array}[]{cc}0&I\\ 0&-\frac{C_{i}}{D_{i}}I\end{array}\right],~~B_{i}=\left[\begin{array}[]{c}0\\ \frac{1}{D_{i}}I\end{array}\right].$
Let the parameters in cost function (49) be given by
$Q_{11}=Q_{22}=Q_{33}=Q_{44}=Q_{55}=0,$
$Q_{12}=Q_{15}=Q_{23}=Q_{34}=Q_{45}=I,$
$Q_{21}=Q_{51}=Q_{32}=Q_{43}=Q_{54}=I,$
$Q_{13}=Q_{14}=Q_{24}=Q_{25}=Q_{35}=0,$
$Q_{31}=Q_{41}=Q_{42}=Q_{52}=Q_{53}=0,$
$R_{1}=R_{2}=R_{3}=R_{4}=R_{5}=I.$
In particular, the initial positions and velocities of the UGVs are given by
$x_{1}(0)=\left[\begin{array}[]{cccc}2\\ 6\\ 1\\ 1\end{array}\right],x_{2}(0)=\left[\begin{array}[]{cccc}3\\ 3\\ 2\\ 2\end{array}\right],x_{3}(0)=\left[\begin{array}[]{cccc}2\\ 2\\ 1\\ 2\end{array}\right],$
$x_{4}(0)=\left[\begin{array}[]{cccc}1\\ 2\\ 1.5\\ 1.5\end{array}\right],x_{5}(0)=\left[\begin{array}[]{cccc}1\\ 4\\ 2\\ 1\end{array}\right].$
By selecting step size $\alpha_{k}=1/k$, $\gamma=2.5$, $\varepsilon=\varepsilon_{1}=\varepsilon_{2}=10^{-4}$, we execute the distributed iterative equations (55)-(58). Then it follows that:
  • The distributed solutions $\|P_{i,\bar{k}}^{\bar{n}}\|_{2}$ with $\bar{n}=20$, $\bar{k}=900$ are:
    $\|P_{1,\bar{k}}^{\bar{n}}\|_{2}=33.5490,~\|P_{2,\bar{k}}^{\bar{n}}\|_{2}=33.5134,$
    $\|P_{3,\bar{k}}^{\bar{n}}\|_{2}=33.5116,~\|P_{4,\bar{k}}^{\bar{\bar{n}}}\|_{2}=33.5111,$
    $\|P_{5,\bar{k}}^{n}\|_{2}=33.5475,$
    which approximate the centralized solution $\|P\|_{2}=33.5591$ solved by (53).
  • The trajectories of $x_{i,\bar{k}}^{\diamond}(t)$ for all UGVs are shown in Fig. 9. It is obvious that $x_{i,k}^{\diamond}(t)$ converges to the centralized solution $x^{\diamond}(t)$ solved by (54).
  • The trajectory of $u(t)$, composed of all distributed controllers $u_{i}(t)$, is shown in Fig. 10, which demonstrates that $u(t)$ converges to the centralized optimal controller $u^{\diamond}(t)$ designed in (52).
  • The position and velocity trajectories of the five UGVs are shown in Figs. 11 and 12, respectively, which demonstrate that all UGVs achieve position consensus, and their velocities eventually converge to zero.
In particular, we conduct a comparative simulation with the widely used classical consensus protocol [38] under the same initial conditions and system parameters $C_{i}$, $D_{i}$. The results are shown in Table II, from which it is obvious that the performance index $J$ of the proposed method is smaller than that obtained by using the controllers in [38].
TABLE II: Performance index comparison between the proposed method and [38] under different $(C_{i},D_{i})$.
Case $C_{i},D_{i}$ $J$ (Proposed Method) $J$ ([38])
Case 1 (0.6, 5) 82.7573 125.1136
Case 2 (0.8, 4) 76.2762 107.3754
Case 3 (1.2, 6) 89.3999 135.6383
Case 4 (1, 4) 76.4114 106.4391
Case 5 (0.4, 5) 82.7004 127.5681
Fig. 9. The trajectories of $\|x_{i,\bar{k}}^{\diamond}(t)\|_{2}$.
Fig. 10. The trajectories of $\|u(t)\|_{2}$.
Fig. 11. The trajectories of the position states $p_{i}(t)$ for all UGVs.
Fig. 12. The trajectories of the velocity states $v_{i}(t)$ for all UGVs.
VII Conclusions